# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v2

@@ Abstract
We present the first application of quantics tensor trains (QTTs) and tensor cross interpolation (TCI) to the solution of a full set of self-consistent equations for multivariate functions, the so-called parquet equations. We show that the steps needed to evaluate the equations (Bethe–Salpeter equations, parquet equation and Schwinger–Dyson equation) can be decomposed into basic operations on the QTT-TCI (QTCI) compressed objects. The repeated application of these operations does not lead to a loss of accuracy beyond a specified tolerance and the iterative scheme converges even for numerically demanding parameters. As examples we take the Hubbard model in the atomic limit and the single impurity Anderson model, where the basic objects in parquet equations, the two-particle vertices, depend on three frequencies, but not on momenta. The results show that this approach is able to overcome major computational bottlenecks of standard numerical methods. The applied methods allow for an exponential increase of the number of grid points included in the calculations leading to an exponentially improving computational error for a linear increase in computational cost.
@@ I Introduction
The understanding of many important excitations of electronic systems – magnons, excitons, or other composite objects – requires understanding correlations at the two particle level. Two-particle quantities – correlation functions or scattering amplitudes (vertices) – are inherently large objects, with multiple dependencies: If we consider scattering of two particles, the amplitude will depend on the energies, momenta and spin-orbitals of two incoming and two outgoing particles. The number of independent variables can be reduced using conservation laws, but each independent spin-orbital combination still depends on three momenta and three frequencies. Numerical representation of these multivariate functions on uniform grids is very expensive due to the third power scaling of memory in the number of discrete momenta or energies. On the other hand, large ranges are required to faithfully represent complicated structures which the vertices show in all their dependencies [1, 2, 3]. When the vertices are themselves variables in diagrammatic equations, as it is the case in parquet equations [4, 5, 6], the required computation time becomes prohibitive [7]. Several solutions to this problem have been proposed so far, either based on partial reduction of the number of frequency and/or momentum variables that need to be treated on grids [8, 9, 10, 11, 12] or based on compact representation of the frequency dependence in a suitable basis [13, 14, 15, 16]. The former still do not lead to true dimensional reduction of the full parquet equations problem. The latter are very promising and provide another path to dimensional reduction, alternative to the one described in this work.
@@
The quantics tensor train (QTT) representation of multivariate functions has already been around for a decade or so [17, 18, 19, 20], but it was only recently applied to various fields of natural science such as turbulence [21, 22, 23, 24, 25], plasma physics [26], quantum chemistry [27], and quantum field theory of the many electron problem [28]. For quantum field theories, the QTT representation provides a compact representation of the space-time dependence of the correlation functions [28]. First many-body calculations of Feynman diagrams with the QTT representation in imaginary time [29] and in nonequilibrium [30] already show the potential of the method. A very favorable scaling of the QTT representation with temperature has been conjectured in Ref. 31. In parallel, the tensor cross interpolation (TCI) method was applied to evaluations of diagrams in many-body physics [32, 33, 34, 35]. TCI can be combined with the quantics tensor train representation to form QTT+TCI=QTCI, a powerful approach with diverse applications [36].
@@ II Models
In the current work we focus on the solution of equations for two-particle vertices in the (Matsubara) frequency space. Although in general the vertices are also dependent on momentum and orbital degrees of freedom, we limit ourselves to simple models for which the vertices depend only on frequency but not on momentum. We present results for two benchmark models: the Hubbard atom, where exact analytical expressions for the vertex functions are known [37] and the single impurity Anderson model [38], where high-quality numerical data is available  [3]. The treatment of the frequency dependence of vertices presented in this work can be directly extended to models with additional orbital and momentum dependencies.
@@ III Parquet equations
The parquet equations are a set of exact relations between different classes of two-particle vertices and between the self-energy and the full two-particle vertex. Before we introduce the equations themselves, we first recapitulate some definitions in order to set the notation.
@@ III.6 Iterative parquet scheme
Assuming we know the fully irreducible vertex $\Lambda$ (or have a good approximation for it) the parquet equation (15) together with the BSEs (12) and the SDE (17), as well as the Dyson equation (6), applied iteratively, will generate all the vertices for the model given by the Hamiltonian and $G_{0}$: {$F_{d/m}$, $F_{s/t}$, $\Gamma_{d/m}$, $\Gamma_{s/t}$, $\Phi_{d/m}$, $\Phi_{s/t}$} as well as the self-energy $\Sigma$ and the Green’s function $G$. The solution is obtained once the iterative scheme converges to a physical solution.
@@ IV.1 Quantics tensor train representation
In the quantics representation, a discrete function $f(m)$ with $m\in\{0,\ldots,M-1\}$ on a one-dimensional grid with $M=2^{R}$ grid points is instead seen as a $2\times 2\times...\times 2$ tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ (see Fig. 2), where each tensor index $\sigma_{1},\ldots,\sigma_{R}$ corresponds to a bit in a binary representation of $m$:
$m=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},\quad\sigma_{\ell}\in\{0,1\}\,.$ (19)
Now, each bit corresponds to a distinct length scale of the system. The first bit $\sigma_{1}$ represents the coarsest length scale which splits the system in halves, while the last bit $\sigma_{R}$ reflects the finest length scale.
@@ IV.2 Tensor Cross Interpolation (TCI)
TCI is more efficient than the SVD-based factorization, especially when the full tensor does not fit into the available memory [32, 36, 34]. SVD-based factorization requires reading all elements of the tensor, leading to an exponential growth of the computation time in $R$. In contrast, if the target tensor/function is low-rank, the computation time of the TCI-based factorization is linear in $R$, leading to an exponential speed-up over the SVD [36]. We refer the reader to Refs. 36, 34 for more technical details.
@@ IV.3 Matrix product operator (MPO)
As we will see in later sections, many operations in QTT can be implemented as the contraction of two MPOs, illustrated in Fig. 3(b). The exact contraction will result in an MPS of large bond dimension $D_{\mathrm{A}}D_{\mathrm{B}}$, where $D_{\mathrm{A}}$ and $D_{\mathrm{B}}$ are the bond dimensions of the two input MPOs, respectively. Thus, the bond dimension of the resulting MPO must be truncated to some $D_{\mathrm{max}}$. The computational cost of a naive contraction followed by truncation scales $\mathcal{O}(D_{\mathrm{A}}^{3}D_{\mathrm{B}}^{3})$.
@@
However, a more efficient scheme is necessary for case (b). We use two algorithms: the “fit algorithm” fits a new MPOs to the MPO-MPO contraction [41], and the “zip-up algorithm” combines contraction of core tensors with truncation of the bond dimensions [42]. We typically combine these, using the zip-up algorithm to generate an initial guess for the fit algorithm. If the resulting MPO is truncated to bond dimension $D_{\mathrm{max}}=D_{\mathrm{A}}=D_{\mathrm{B}}$, the computational cost of both algorithms scales as $\mathcal{O}(D_{\mathrm{max}}^{4}L)$.
@@ V.1 Quantics representation and compression of two-particle vertex functions
The first step in using the QTT framework for solving the parquet equations is to investigate the compressibility of vertex functions in the above representation. We use the full vertex in the density channel ($F_{d}$) of the Hubbard atom for numerical demonstration. The fermionic frequency dependence of $F_{d}$ at $\omega=0$ is shown for various temperatures in Fig. 5. Note, that in the Hubbard atom the results are not separately dependent on temperature and $U$, but only on their ratio $\beta U$, since there are no other energy scales [37]. Before we focus on the scaling of bond dimension with temperature, let us first look at the bond dimension along the QTT at $\beta U=1$ for different grid sizes and tolerances set in TCI as shown in Fig. 6. Moving inward from the first and last bonds, the bond dimension grows exponentially as $D_{\ell}=\min(2^{\ell},2^{\mathcal{L}-\ell})$, which is the maximum bond dimension of an uncompressed factorization and represents maximum entanglement between these exponentially different length scales. In between, the bond dimension then saturates at a maximum bond dimension $D_{\mathrm{max}}$, therefore indicating that the vertex structures are indeed compressible. The maximum bond dimension $D_{\mathrm{max}}$ is between $80$ and $400$, and increases with decreasing tolerance $\epsilon$. Importantly, $D_{\mathrm{max}}$ is nearly independent of the grid parameter $R$, an exponential increase in the number of grid points ($\mathcal{O}(2^{R})$) can be achieved for linear cost ($\mathcal{O}(R)$) in runtime and memory. These findings are consistent with earlier results on the compressibility of vertices in Ref. 28.
@@
In Fig. 7, we show $D_{\mathrm{max}}$ of the QTT representation of $F_{d}$ for various values of $\beta U$, different tolerances set in TCI, and various grid sizes. As expected from earlier considerations, the maximum bond dimension is quite low at high temperatures and steeply grows with $\beta U$ until $\beta U\approx 5.1$, where it reaches a maximum. Thereafter, the bond dimension decreases again, which does not conform with our expectations. The maximum can be related to the first global divergence of the irreducible vertex $\Gamma_{d}$ [37, 43, 44, 45, 46], which occurs at $\beta U=5.13715$ and is indicated by a dashed line in the plot. Although the irreducible vertex $\Gamma_{d}$ diverges at this value of $\beta U$, there is no phase transition connected with the divergence and $F_{d}$ remains finite. The presence of a maximum in $D_{\mathrm{max}}$ precisely at the first global divergence of $\Gamma_{d}$ is very interesting from two different perspectives. Firstly, we see that $D_{\mathrm{max}}$ stops growing with $\beta U$ and, thus, calculations for temperatures ranging from very low to very high are manageable within the QTT framework in this case. Secondly, although the full vertex $F_{d}$ does not contain singularities, we see fingerprints of this first global divergence of $\Gamma_{d}$ in the amount of length scale entanglement in the system represented by the maximum bond dimension. A deeper investigation of this behavior is left for future work.
@@ V.2 Channel transformations
In this process, there are two distinct sources of error: the finite frequency box and the QTT truncation. As an example, Fig. 9(a) shows the absolute normalized error ($||\Delta F_{\textit{pp}}||:=|F_{\textit{pp},\mathrm{trafo}}-F_{\textit{pp},\mathrm{exact}}|/||F_{\textit{pp},\mathrm{exact}}||_{\infty}$) of a transformation of the full vertex $F$ in the ph channel parametrization to the pp channel parametrization (denoted as $F_{\textit{pp}}$) in the $\nu,\nu^{\prime}$-plane. In two triangular regions, the upper right and lower left corner, errors are large due to the finite size of the frequency box, and are not caused by the QTT compression. These points correspond to grid points outside of the original frequency box that were transformed into the box by the ph to pp transformation. The missing data there can be either replaced by zeros (which corresponds to open boundary conditions of the affine transformation, see App. B) or by values extrapolated from another part of the frequency box (e.g. by using periodic boundary conditions for the affine transformation, as in Ref. [47, footnote 14]). We checked that for the examples shown in this paper the average difference in error between the two options is small. We used periodic boundary conditions for all results presented in the manuscript. For the half-filled Hubbard atom this leads to better representation of missing values on the diagonal due to high symmetry of the full vertex in this case. In general, the choice of boundary conditions can be adapted to the problem at hand. Since the error can easily be reduced and shifted to higher Matsubara frequencies by increasing the size of the frequency box exponentially through increasing $R$, we do not expect the choice of boundary conditions to have significant effect on the error.
In the remaining diagonal region in between the two yellow corners in Fig. 9(a), the error is entirely due to the tensor train approximation. At a bond dimension of $D_{\mathrm{max}}=100$, the error is smaller than the error tolerance of $\epsilon=10^{-5}$ everywhere.

### v3

# Two-particle calculations with quantics tensor trains: Solving the parquet equations

@@ Abstract
We present the first application of quantics tensor trains (QTTs) and tensor cross interpolation (TCI) to the solution of a full set of self-consistent equations for multivariate functions, the so-called parquet equations. We show that the steps needed to evaluate the equations (Bethe–Salpeter equations, parquet equation and Schwinger–Dyson equation) can be decomposed into basic operations on the QTT-TCI (QTCI) compressed objects. The repeated application of these operations does not lead to a loss of accuracy beyond a specified tolerance and the iterative scheme converges even for numerically demanding parameters. As examples we take the Hubbard model in the atomic limit and the single impurity Anderson model, where the basic objects in parquet equations, the two-particle vertices, depend on three frequencies, but not on momenta. The results show that this approach is able to overcome major computational bottlenecks of standard numerical methods. The applied methods allow for an exponential increase of the number of grid points included in the calculations, and a corresponding exponential reduction of the computational error, for a linear increase in computational cost.
@@ I Introduction
The understanding of many important excitations of electronic systems – magnons, excitons, or other composite objects – requires understanding correlations at the two particle level. Two-particle quantities – correlation functions or scattering amplitudes (vertices) – are inherently large objects, with multiple dependencies: If we consider scattering of two particles, the amplitude will depend on the energies, momenta and spin-orbitals of two incoming and two outgoing particles. The number of independent variables can be reduced using conservation laws, but each independent spin-orbital combination still depends on three momenta and three frequencies. Numerical representation of these multivariate functions on uniform grids is very expensive due to the third power scaling of memory in the number of discrete momenta or energies. On the other hand, large ranges are required to faithfully represent complicated structures which the vertices show in all their dependencies [1, 2, 3]. When the vertices are themselves variables in diagrammatic equations, as it is the case in parquet equations [4, 5, 6], the required computation time becomes prohibitive [7]. Several solutions to this problem have been proposed so far, either based on partial reduction of the number of frequency and/or momentum variables that need to be treated on grids [8, 9, 10, 11, 12] or based on compact representation of the frequency dependence in a suitable basis [13, 14, 15, 16]. The former still do not lead to true dimensional reduction of the full parquet equations problem. The latter are very promising and provide another path to dimensional reduction, alternative to the one described in this work. Recently, also a wavelet-based decomposition for efficiently compressing two-particle quantities has been proposed [17, 18, 19].
@@
The quantics tensor train (QTT) representation of multivariate functions has already been around for a decade or so [20, 21, 22, 23], but it was only recently applied to various fields of natural science such as turbulence [24, 25, 26, 27, 28], plasma physics [29], quantum chemistry [30], and quantum field theory of the many electron problem [31]. For quantum field theories, the QTT representation provides a compact representation of the space-time dependence of the correlation functions [31]. First many-body calculations of Feynman diagrams with the QTT representation in imaginary time [32] and in nonequilibrium [33] already show the potential of the method. A very favorable scaling of the QTT representation with temperature has been conjectured in Ref. 34. In parallel, the tensor cross interpolation (TCI) method was applied to evaluations of diagrams in many-body physics [35, 36, 37, 38]. TCI can be combined with the quantics tensor train representation to form QTT+TCI=QTCI, a powerful approach with diverse applications [39].
@@ II Models
In the current work we focus on the solution of equations for two-particle vertices in the (Matsubara) frequency space. Although in general the vertices are also dependent on momentum and orbital degrees of freedom, we limit ourselves to simple models for which the vertices depend only on frequency but not on momentum. We present results for two benchmark models: the Hubbard atom, where exact analytical expressions for the vertex functions are known [40] and the single impurity Anderson model [41], where high-quality numerical data is available  [3]. The treatment of the frequency dependence of vertices presented in this work can be directly extended to models with additional orbital and momentum dependencies. The possibility of such extensions will be discussed in App. G.
@@ III Parquet equations
The parquet equations are a set of exact relations between different classes of two-particle vertices and between the self-energy and the full two-particle vertex [4, 5]. A good introduction to the formalism is provided in [6]. Here we only recall the equations, using the notation of Refs. [1, 42, 43, 44]. Before we introduce the equations themselves, we first recapitulate some definitions in order to set the notation.
@@ III.6 Iterative parquet scheme
Assuming we know the fully irreducible vertex $\Lambda$ (or have a good approximation for it) the parquet equation (15) together with the BSEs (12) and the SDE (17), as well as the Dyson equation (6), applied iteratively, will generate all the vertices for the model given by the Hamiltonian and $G_{0}$: {$F_{d/m}$, $F_{s/t}$, $\Gamma_{d/m}$, $\Gamma_{s/t}$, $\Phi_{d/m}$, $\Phi_{s/t}$} as well as the self-energy $\Sigma$ and the Green’s function $G$. The iterations are repeated until the difference between consecutive values falls below a given tolerance.
@@
We conclude this section with some comments on convergence issues. First, the iterative parquet scheme is not guaranteed to converge. Second, there are cases where the iterative scheme leads to a false solution. This is the case, e.g., for the Hubbard atom for large $\beta U$ beyond the first divergence of the irreducible vertex [45]. In this work we do not address such cases and focus on examples and parameter regimes where the parquet scheme should converge to the physical solution. Nevertheless, convergence can also be influenced by other factors. The first numerical solution for the Hubbard model on a $4\times 4$ cluster [7] showed the difficulty of achieving convergence, particularly when the crossing symmetry (see App. A) was not obeyed. An important factor in improving stability of the iterative scheme turned out to be the inclusion of vertex asymptotics, i.e., a prediction for values of the vertex that fall beyond the frequency range used, either by introducing so-called kernels [46, 44, 10, 8], or by removing the asymptotic parts of vertices altogether in the single-boson exchange reformulation [12, 3]. In this work we do not use asymptotics and do not suffer from convergence problems mainly because we are able to use very large grids. In the future it might be important to include asymptotic, as in Refs. [46, 44, 10, 8] or [12, 3], to improve stability but also to avoid some technical issues that are addressed in App. F.
@@ IV Quantics Tensor trains
In the remainder of the paper, the so-called grid parameter $R$ will be of central importance, where the three-dimensional Matsubara frequency grid will consist of $2^{3R}$ grid points. Hence, this parameter governs the (exponential) number of points of the discretized grid and exponentially different length scales in the system and, thus, determines the length of the resulting QTT. This will become more clear in the following sections.
@@ IV.1 Quantics tensor train representation
Before discussing the three-dimensional Matsubara frequency case, let us introduce the QTT formalism for the one-dimensional case for educational purposes. In the quantics representation, a discrete function $f(m)$ with $m\in\{0,\ldots,M-1\}$ on a one-dimensional grid with $M=2^{R}$ grid points is instead seen as a $2\times 2\times...\times 2$ ($R$ times) tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ (see Fig. 2), where each tensor index $\sigma_{1},\ldots,\sigma_{R}$ corresponds to a bit in a binary representation of $m$:
$m=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},\quad\sigma_{\ell}\in\{0,1\}\,,$ (19)
with the discussed grid parameter $R$. Now, each bit corresponds to a distinct length scale of the system. The first bit $\sigma_{1}$ represents the coarsest length scale which splits the system in halves, while the last bit $\sigma_{R}$ reflects the finest length scale. Hence, for continuous variables $m$ defined on a specific interval the grid parameter $R$ determines how exponentially dense the grid gets, while for discrete variables like Matsubara frequencies, $R$ determines how exponentially large the grid gets. In both cases, $R$ specifies the (exponential) number of grid points and a linear increase in $R$ corresponds to an exponential increase in the number of grid points.
@@ IV.2 Tensor Cross Interpolation (TCI)
TCI is more efficient than the SVD-based factorization, especially when the full tensor does not fit into the available memory [35, 39, 37]. SVD-based factorization requires reading all elements of the tensor, leading to an exponential growth of the computation time in $R$. In contrast, if the target tensor/function is low-rank, the computation time of the TCI-based factorization is linear in $R$, leading to an exponential speed-up over the SVD [39]. We refer the reader to Refs. 39, 37 for more technical details , e.g. for information on how the sampling points in the TCI algorithm are chosen. In the following computations TCI will only be used for compressing the initial input vertices and functions. On a more technical note, the crossinterpolate2 algorithm in the TensorCrossInterpolation.jl library was used for compressing the objects. More details on this specific algorithm can be found in section 8.3.1 in Ref. 37.
@@ IV.3 Matrix product operator (MPO)
As we will see in later sections, many operations in QTT can be implemented as the contraction of two MPOs, illustrated in Fig. 3(b). The exact contraction will result in an MPS of large bond dimension $D_{\mathrm{A}}D_{\mathrm{B}}$, where $D_{\mathrm{A}}$ and $D_{\mathrm{B}}$ are the bond dimensions of the two input MPOs, respectively. Thus, the bond dimension of the resulting MPO must be truncated to some $D_{\mathrm{max}}$. The computational cost of a naive SVD-based contraction followed by truncation scales $\mathcal{O}(D_{\mathrm{A}}^{3}D_{\mathrm{B}}^{3})$.
@@
However, a more efficient scheme is necessary for case (b). We use two algorithms: the “fit algorithm” fits a new MPOs to the MPO-MPO contraction [48], and the “zip-up algorithm” combines contraction of core tensors with truncation of the bond dimensions [49]. We typically combine these, using the zip-up algorithm to generate an initial guess for the fit algorithm. If the resulting MPO is truncated to bond dimension $D_{\mathrm{max}}=D_{\mathrm{A}}=D_{\mathrm{B}}$, the computational cost of both algorithms scales as $\mathcal{O}(D_{\mathrm{max}}^{4}L)$. We want to emphasize that in the current implementation the combination of the two algorithms is still SVD- and not TCI-based. Thus, TCI is only used in the compression of the input functions and not in the MPO-MPO contractions. However, in future work we plan on using TCI for MPO-MPO contractions as well, since this is expected to be computationally more efficient.
@@ V.1 Quantics representation and compression of two-particle vertex functions
The first step in using the QTT framework for solving the parquet equations is to investigate the compressibility of vertex functions in the above representation. We use the full vertex in the density channel ($F_{d}$) of the Hubbard atom for numerical demonstration. The fermionic frequency dependence of $F_{d}$ at $\omega=0$ is shown for various temperatures in Fig. 5. Note, that in the Hubbard atom the results are not separately dependent on temperature and $U$, but only on their ratio $\beta U$, since there are no other energy scales [40]. Before we focus on the scaling of bond dimension with temperature, let us first look at the bond dimension along the QTT at $\beta U=1$ for different grid sizes and tolerances set in TCI as shown in Fig. 6. Moving inward from the first and last bonds, the bond dimension grows exponentially as $D_{\ell}=\min(2^{\ell},2^{\mathcal{L}-\ell})$, which is the maximum bond dimension of an uncompressed factorization and represents maximum entanglement between these exponentially different length scales. In between, the bond dimension then saturates at a maximum bond dimension $D_{\mathrm{max}}$, therefore indicating that the vertex structures are indeed compressible. The maximum bond dimension $D_{\mathrm{max}}$ is between $80$ and $400$, and increases with decreasing tolerance $\epsilon$. Importantly, $D_{\mathrm{max}}$ is nearly independent of the grid parameter $R$, an exponential increase in the number of grid points ($\mathcal{O}(2^{R})$) can be achieved for linear cost ($\mathcal{O}(R)$) in runtime and memory. These findings are consistent with earlier results on the compressibility of vertices in Ref. 31.
@@
In Fig. 7, we show $D_{\mathrm{max}}$ of the QTT representation of $F_{d}$ for various values of $\beta U$, different tolerances set in TCI, and various grid sizes. As expected from earlier considerations, the maximum bond dimension is quite low at high temperatures and steeply grows with $\beta U$ until $\beta U\approx 5.1$, where it reaches a maximum. Thereafter, the bond dimension decreases again, which does not conform with our expectations. The maximum can be related to the first global divergence of the irreducible vertex $\Gamma_{d}$ [40, 50, 51, 52, 53, 54], which occurs at $\beta U=5.13715$ and is indicated by a dashed line in the plot. Although the irreducible vertex $\Gamma_{d}$ diverges at this value of $\beta U$, there is no phase transition connected with the divergence and $F_{d}$ remains finite. The presence of a maximum in $D_{\mathrm{max}}$ precisely at the first global divergence of $\Gamma_{d}$ is very interesting from two different perspectives. Firstly, we see that $D_{\mathrm{max}}$ stops growing with $\beta U$ and, thus, calculations for temperatures ranging from very low to very high are manageable within the QTT framework in this case. Secondly, although the full vertex $F_{d}$ does not contain singularities, we see fingerprints of this first global divergence of $\Gamma_{d}$ in the amount of length scale entanglement in the system represented by the maximum bond dimension. A deeper investigation of this behavior is left for future work.
@@
The Hubbard atom is quite an extreme limit of the Hubbard model and one could think that the high compressibility is related to structures present in the vertices that are limited to this atomic limit. However, this is not the case: prominent structures in the frequency dependence of vertices are well-known to be present in the SIAM (see Sec. VI below for an example), as well as in the local approximation of the Hubbard model (in dynamical mean-field theory) [42, 52]. These structures arise specifically from certain types of diagrams [42, 52] and also from insertions from one scattering channel to another through the parquet equation (15). For models whose vertices depend on momentum and/or orbital indices, the structures may become more complicated. Nevertheless, their origin is well understood and we expect the vertices to generically have structures for a large range of parameters and models. As long as these structures show some kind of scale separation, they will be QTT compressible, hopefully with still manageable bond dimensions.
@@ V.2 Channel transformations
In this process, there are two distinct sources of error: the finite frequency box and the QTT truncation. As an example, Fig. 9(a) shows the absolute normalized error ($||\Delta F_{\textit{pp}}||:=|F_{\textit{pp},\mathrm{trafo}}-F_{\textit{pp},\mathrm{exact}}|/||F_{\textit{pp},\mathrm{exact}}||_{\infty}$) of a transformation of the full vertex $F$ in the ph channel parametrization to the pp channel parametrization (denoted as $F_{\textit{pp}}$) in the $\nu,\nu^{\prime}$-plane. In two triangular regions, the upper right and lower left corner, errors are large due to the finite size of the frequency box, and are not caused by the QTT compression. These points correspond to grid points outside of the original frequency box that were transformed into the box by the ph to pp transformation. The missing data there can be either replaced by zeros (which corresponds to open boundary conditions of the affine transformation, see App. B) or by values extrapolated from another part of the frequency box (e.g. by using periodic boundary conditions for the affine transformation, as in Ref. [55, footnote 14]). We checked that for the examples shown in this paper the average difference in error between the two options is small. We used periodic boundary conditions for all results presented in the manuscript. For the half-filled Hubbard atom this leads to better representation of missing values on the diagonal due to high symmetry of the full vertex in this case. In general, the choice of boundary conditions can be adapted to the problem at hand. Since the error can easily be reduced and shifted to higher Matsubara frequencies by increasing the size of the frequency box exponentially through increasing $R$, we do not expect the choice of boundary conditions to have significant effect on the error.
In the remaining diagonal region in between the two yellow corners in Fig. 9(a), the error is entirely due to the tensor train approximation. At a bond dimension of $D_{\mathrm{max}}=100$, the error is smaller than the error tolerance of $\epsilon=10^{-5}$ everywhere, meaning that the crossing symmetry is also fulfilled up to this error.